Amber rolls a 6-sided die. On her first roll, she gets a "4". She rolls again.
(a) What is the probability that the second roll is also a "4".
P(4 | 4) =
(b) What is the probability that the second roll is a "2".
P(2 / 4) =

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3835)


The answer to each question is 1/6 since there is 1 side desired out of 6 sides total.

It does not matter what the previous roll was. Each roll of the die is independent from any other.

Answer by ikleyn(53763)

.
Amber rolls a 6-sided die. On her first roll, she gets a "4". She rolls again.
(a) What is the probability that the second roll is also a "4".
P(4 | 4) =
(b) What is the probability that the second roll is a "2".
P(2 / 4) =
~~~~~~~~~~~~~~~~~~~~~~~~~~~


For both your questions, (a) and (b), the outcomes of the second roll are independent from outcomes of the first roll.

THEREFORE,  

    in question  (a)   P(4 | 4) = 1/6,  the same as  P(4)  in the first roll and/or  P(4) in the second roll.

         To write here P(4 | 4)  is somehow awkward, if you don't want to deceive yourself and people around, 
         because the outcomes of the first roll and of the second roll are independent
         and this independence is obvious and is well known to all, to anyone and to everybody.



    in question (b)  P(2 | 4) = 1/6,  the same as  P(2)  in the first roll and/or  P(2) in the second roll.

         To write here P(2 | 4)  is somehow awkward, if you don't want to deceive yourself and people around, 
         because the outcomes of the first roll and the second roll are independent
         and this independence is obvious and is well known to all, to anyone and to everybody.